COMBINATION BANDS BETWEEN 2900 AND 3600 CM^{−1} OF CYCLIC O_{4} CATION TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Sensor Science Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

The infrared spectrum of cyc-O_{4}^{+} trapped in solid neon includes a group of combination bands built on (ν_{1} + ν_{5}) of ground-state cyc-O_{4}^{+}. Each peak lies close to a counterpart previously reported^{a} in a study of the infrared laser photodissociation spectroscopy of a mass-selected molecular beam. This agreement is consistent with the vibrational assignment of three low-frequency modes of cyc-O_{4}^{+} that was proposed in the earlier study. The spectra obtained for the isotopologues formed by the substitution of one or two ^{18}O_{2} moieties suggest the occurrence of interaction with a nearby excited electronic state.

aA. M. Ricks, G. E. Douberly, and M. A. Duncan, Int. J. Mass Spectrom. 283, 69 (2009).